OpenStudy (anonymous):

integral from x to (x^2) of (t^4)dt

zepdrix (zepdrix):

\[\large \int\limits_x^{x^4}t^4\;dt\] Simply apply the `Power Rule for Integration`. ~Increase the power on t by 1. ~Divide by this new power.

zepdrix (zepdrix):

Can you figure that part out? After that, we have to evaluate the function at the upper and lower limits of integration. :)

OpenStudy (anonymous):

doesnt the t^4 become (t^5)/5

zepdrix (zepdrix):

Yes good.

OpenStudy (anonymous):

i am going to make a guess that you are asked to find the derivative of that thing i may be wrong

zepdrix (zepdrix):

Yah it looks like an FTC problem :p But whatev :D

zepdrix (zepdrix):

\[\large \color{royalblue}{\frac{1}{5}t^5}|_x^{x^4}\]So here the anti-derivative of our function we started with. We now will plug in our upper boundary `x^4` and subtract from that, the function evaluated at the lower limit.\[\large \color{royalblue}{\frac{1}{5}t^5}|_x^{x^4} \quad = \quad \left(\color{royalblue}{\frac{1}{5}(x^4)^5}\right)-\left(\color{royalblue}{\frac{1}{5}(x)^5}\right)\]Understand what we did there Umbaka? :O

OpenStudy (anonymous):

thanks, I know how to get the antiderivative but since it is asking for the derivative, does it become [(4(x^6))(2x)]-[4x^3]

zepdrix (zepdrix):

Hmmm I think the derivative will give us this, \[\large \frac{d}{dx}\left(\color{royalblue}{\frac{1}{5}x^{20}}\right)-\left(\color{royalblue}{\frac{1}{5}x^5}\right)\] \[\large =4x^{19}-x^4\]

OpenStudy (anonymous):

Thank you so much!

OpenStudy (sirm3d):

aren't you suppose to use the fundamental theorem of calculus?

zepdrix (zepdrix):

Yes, but he didn't inform us that we were taking a derivative until we had already gone through most of the problem :) So we just finished it this way.